# Approximating Sparse Matrices and their Functions using Matrix-vector products

### v2

@@ 2 Approximation for functions of general sparse matrices
Let $\bm{A}$ be a diagonalizable sparse matrix with some sparsity pattern. Suppose that $f$ is analytic in a region containing the spectrum of $\bm{A}$. Then depending on the sparsity pattern of $\bm{A}$, the entries of $f(\bm{A})$ decay exponentially away from certain small regions of the matrix. More formally, associate the unweighted directed graph $G(\bm{A})=(V,E)$ to $\bm{A}=[A_{ij}]$ so that $\bm{A}$ is the adjacency matrix of $G(\bm{A})$. $V$ is the vertex set consisting of integers from $1$ to $n$ and $E$ is the edge set containing all ordered pairs $(i,j)$ with $A_{ij}\neq 0$. We define the distance $d_{G}(i,j)$ between vertex $i$ and vertex $j$ to be the shortest directed path from $i$ to $j$ [^2] Note that if $\bm{A}$ is not symmetric then $d_{G}(i,j)\neq d_{G}(j,i)$ in general. We then have that there exist constants $C_{0}>0$ and $10<\lambda_{0}<$ 1 such that
@@
for all $i,j=1,2,...,n$. For further details see [8]. This result tells us that if $d_{G}(i,j)\gg 1$ then $|[f(\bm{A})]_{ij}|\ll 1$, which in turn means that depending on the sparsity pattern of $\bm{A}$, there may only be few large entries in $f(\bm{A})$ with the rest being negligible. With this observation we can aim to recover the dominating entries of each row/column of $f(\bm{A})$. We solve this problem using compressed sensing by interpreting small entries as noise and the large entries as signals. [^3] Another possibility would be to identify the nonzero structure using the graph structure $(V,E)$ and $f$. This can quickly become intractable, and in some cases $f(\bm{A})$ can be sparse despite the graph structure suggesting otherwise, so here we opt for a generic approach.
Compressed sensing [13, 23] is a technique in signal processing for efficiently reconstructing a sparse signal. Let $\bm{Y}\in\mathbb{R}^{n\times s}$ be a so-called sensing operator and $\bm{v}\in\mathbb{R}^{n}$ be an unknown sparse vector. Then the problem of recovering $\bm{v}$ using the measurements $\bm{y}=\bm{Y}^{T}\bm{v}$ is a combinatorial optimization problem given by
@@
where $\left\lVert\bm{z}\right\rVert_{0}$ denotes the number of nonzeros in the vector $\bm{z}$. There are two established classes of algorithms, among others, to solve (2) in the compressed sensing literature, namely basis pursuit or $\ell_{1}$-minimization using convex optimization [14], and greedy methods [12]. For large $n$, convex optimization can become infeasible (especially given that we need to solve (2) $n$ times to recover the whole matrix), so we use the cheaper option, greedy methods. Greedy methods are iterative and at each iteration attempt to find the correct location and the values of the dominant entries.
Let $\bm{v}\in\mathbb{R}^{n}$ be a $k$-sparse vector, that is, a vector with at most $k$ nonzero entries. To recover $\bm{v}$ with theoretical guarantee we typically require at least $s=\mathcal{O}(k\log n)$ measurements. In practice, $s=\mathcal{O}(k)$ measurements seem to work well [10]. Each measurement is taken through vector-vector multiply with $\bm{v}$. The three classes of frequently used sensing operators $\bm{Y}$ are: a Gaussian matrix $\bm{\mathcal{N}}$ with entries drawn from i.i.d. $\mathcal{N}(0,s^{-1})$, $s$ columns of the discrete cosine transform matrix $\bm{\mathcal{C}}$ sampled uniformly at random and a sparse matrix $\bm{\mathcal{S}}_{\xi}$ with $\xi$ nonzero entries per row with the support set drawn uniformly at random and the nonzero values drawn uniformly at random from ±$\pm\xi^{-1/2}$. Here, $\xi$ is the sparsity parameter for the sparse sensing operator.
There are a number of greedy methods available, including Normalized Iterative Hard Thresholding (NIHT) [11], Hard Thresholding Pursuit (HTP) [27] and Compressive Sampling Matching Pursuit (CoSaMP) [51]. Among the key differences and the trade-offs between the different methods is the cost per iteration in order to speed up support detection and the number of iterations for better asymptotic convergence rate. Many greedy methods have uniform recovery guarantees if the sensing operator $\bm{Y}$ has sufficiently small restricted isometry constants [28]. [^4] Under mild assumptions on $s$, the three classes of sensing operators described in the previous paragraph have small restricted isometry constants with high probability. This means that if we want to recover a matrix then we can use the same sensing operator $\bm{Y}$ for the recovery of all rows of a matrix, which makes it much more efficient than recovering $n$ rows of a matrix individually using different sensing operators.
@@
Here $H_{k}$ is the hard thresholding operator that sends all but the largest $k$ entries (in absolute value) of the input vector to 00. The stepsize $\mu$ in NIHT is chosen to be optimal when the support of the current iterate is the same as the sparsest solution. Assuming that $\bm{Y}$ satisfies the restricted isometry property (RIP), then at iteration $N$, the output $\bm{v}^{(N)}$ of NIHT satisfies
@@
Further details about NIHT can be found in [11].
The dominating cost for the greedy methods is the matrix-vector product between the sensing operator or the transpose of the sensing operator and the unknown sparse signal, i.e., the cost of $\bm{v}\mapsto\bm{Y}\bm{v}$ and $\bm{w}\mapsto\bm{Y}^{T}\bm{w}$ respectively. If the sensing operator is a Gaussian $\bm{\mathcal{N}}$ then the cost is $\mathcal{O}(ns)$, if $\bm{Y}$ is a subsampled discrete cosine matrix $\bm{\mathcal{C}}$ then the cost is $\mathcal{O}(n\log n)$ or $\mathcal{O}(n\log s)$ if we use the subsampled FFT algorithm [55] and lastly, if $\bm{Y}$ is a sparse matrix $\bm{\mathcal{S}}_{\xi}$ then the cost is $\mathcal{O}(\xi n)$.
@@ 2.1 Strategy for recovering general sparse matrices
where $C_{B}>0$ and $10<\lambda_{B}<$ 1 are constants and $d(i,j)\in\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is some bivariate function that captures the magnitude of the entries of $\bm{B}$. Functions of sparse matrices $\bm{B}=f(\bm{A})$ with a smooth $f$ is a special case of $\bm{B}$ that satisfies (1). Another special case is $\bm{B}=\bm{A}$, which recovers the sparse matrix $\bm{A}$ itself. The zero entries of $\bm{B}$ would correspond to $d(i,j)=\infty$ in (5).
[…]
@@
2:$\widehat{\bm{B}}$ sparse matrix with at most $k$ nonzero entries in each row.
3:Draw a sensing operator $\bm{Y}\in\mathbb{R}^{n\times s}$, e.g. Gaussian. (rec. $s=\lceil 2k\log(n/k)\rceil$)
@@
5:for $i=1,2,...,n$ do
6:     $\widehat{\bm{B}}(i,:)=\texttt{CS}(\bm{Y},\bm{F}_{s}(i,:),k)$triangleright Using compressed sensing to approximate the rows of $\bm{B}$
@@
The algorithm SpaMRAM is a straightforward application of compressed sensing techniques used to recover a sparse matrix instead of a vector, particularly when the goal is to recover $\bm{A}$ itself. We are nonetheless unaware of existing work that spells out explicitly, and SpaMRAM can be significantly more efficient than alternative approaches [41, 53] that vectorize the matrix. Since most sensing matrices satisfy the RIP with failure probability exponentially decaying in $s$ [28], by the union bound we can conclude that the whole matrix can be recovered with high probability.
@@
When $\bm{B}=f(\bm{A})$ for sparse $\bm{A}$, we can evaluate $f(\bm{A})\bm{Y}$ using Krylov methods. Assuming that the Krylov method converges (to sufficient accuracy) in $\mathcal{O}(1)$ iterations, the complexity of SpaMRAM is $\mathcal{O}(s\cdot nnz(\bm{A})+n^{2}\log n)=\mathcal{O}(k\log n\cdot nnz(\bm{A})+n^{2}\log n)$ which consists of $\mathcal{O}(s\cdot nnz(\bm{A}))=\mathcal{O}(k\log n\cdot nnz(\bm{A}))$ flops for computing $s=\mathcal{O}(k\log n)$ matrix-vector products with $f(\bm{A})$ and $\mathcal{O}(n^{2}\log n)$ flops for solving $n$ compressed sensing problems. Here $nnz(\bm{A})$ is the number of nonzero entries of $\bm{A}$.
@@ 2.1.2 Dependency of the compressed sensing solver in SpaMRAM
The majority of the computational task in SpaMRAM (Algorithm 1) relies on compressed sensing. There are three main parameter choices in relation to compressed sensing for SpaMRAM: the sensing operator $\bm{Y}$, the number of measurements $s$ and the compressed sensing solver CS. For the solver, as mentioned earlier in Section 2, different greedy algorithms have similar guarantees; see [28] and their relevant papers or for an empirical study see [10]. In this work, we focus on, and suggest NIHT as it is the simplest and has a low cost per iteration. For a greedy method with a better asymptotic guarantee, but with a higher cost per iteration, we suggest CoSaMP or its variants, e.g. CSMPSP discussed in [10]. For the sensing operator $\bm{Y}$, all three choices mentioned previously are good choices. For robustness, we recommend the Gaussian matrix $\bm{\mathcal{N}}$ and for speed, we recommend the sparse matrix $\bm{\mathcal{S}}_{\xi}$ with the sparsity parameter $\xi=\min\{8,s\}$ if a good sparse matrix library is available, otherwise we recommend the subsampled DCT matrix $\bm{\mathcal{C}}$. Lastly, for the number of measurements $s$ we recommend $s=\mathcal{O}\left(k\log\left(\frac{n}{k}\right)\right)$ for theoretical guarantee [23, 13], which is suggested in SpaMRAM with $s=\lceil 2k\log(n/k)\rceil$. However, a constant multiple of the sparsity parameter $k$, say $s=10k$, is observed to work well in practice [10, 28].
@@ 2.1.3 A posteriori error estimate for SpaMRAM
The behaviour of SpaMRAM can be difficult to grasp as it relies on various parameters that are sometimes implicit, e.g., (5). Therefore, we can incorporate a posteriori error testing for SpaMRAM to observe how well it might be perform on a particular problem. Let us try estimate the error using known quantities. We have the sensing operator $\bm{Y}$, the measurements $\bm{F}_{s}\in\mathbb{R}^{n\times s}$ and the output of SpaMRAM, $\hat{\bm{B}}$. If the matrix-vector products (line $2$ of SpaMRAM) have been computed exactly then we have $\bm{B}\bm{Y}=\bm{F}_{s}$, so we try estimate the error by determining how close the measurements computed using the output of SpaMRAM, i.e. $\hat{\bm{B}}\bm{Y}$, is to the computed measurements $\bm{F}_{s}$. We define
@@
as the relative error of the computed solution of SpaMRAM quantified through the measurements. Note that the error estimate $\delta_{RE}^{(S)}$ can be computed a posteriori. The error estimate $\delta_{RE}^{(S)}$ can be used as a heuristic for how SpaMRAM performs. The main motivation for this approach is that a small number of Gaussian measurements estimate the norm well [35, 48]. In our case, we essentially have $\bm{F}_{s}=\bm{B}\bm{Y}$ and $\hat{\bm{B}}\bm{Y}$ giving us $(\bm{B}-\hat{\bm{B}})\bm{Y}$, which is a small number of Gaussian measurements of the absolute error $\bm{B}-\hat{\bm{B}}$ if $\bm{Y}$ is the Gaussian sensing operator. [^7] For theoretical guarantee, $\bm{Y}$ needs to be independent from $\bm{B}$ and $\hat{\bm{B}}$ [35], however the output of SpaMRAM, $\hat{\bm{B}}$ is dependent on $\bm{Y}$. Nonetheless, the dependency is rather weak so $\delta_{RE}^{(S)}$ should be a good error estimate in practice. For theoretical guarantee, one can take, say $5$ extra measurements for error estimation. Therefore $\delta_{RE}$ provides a good guidance for how small the error is for SpaMRAM. For the other sensing operators, the behaviour is similar [48]. Later, we show in a numerical experiment (Figure 2) in Section 4 that $\delta_{RE}^{(S)}$ can estimate the performance of SpaMRAM well.

### v3

@@ 2 Approximation for functions of general sparse matrices
Let $\bm{A}$ be a diagonalizable sparse matrix with some sparsity pattern. Suppose that $f$ is analytic in a region containing the spectrum of $\bm{A}$. Then depending on the sparsity pattern of $\bm{A}$, the entries of $f(\bm{A})$ decay exponentially away from certain small regions of the matrix. More formally, associate the unweighted directed graph $G(\bm{A})=(V,E)$ to $\bm{A}=[A_{ij}]$ so that $\bm{A}$ is the adjacency matrix of $G(\bm{A})$. $V$ is the vertex set consisting of integers from $1$ to $n$ and $E$ is the edge set containing all ordered pairs $(i,j)$ with $A_{ij}\neq 0$. We define the distance $d_{G}(i,j)$ between vertex $i$ and vertex $j$ to be the shortest directed path from $i$ to jj. [^2] Note that if $\bm{A}$ is not symmetric then $d_{G}(i,j)\neq d_{G}(j,i)$ in general. We then have that there exist constants $C_{0}>0$ and $0<\lambda_{0}<1$ such that
@@
for all i,j=1,2,,ni,j=1,2,...,n. For further details see Benzi2007. This result tells us that if $d_{G}(i,j)\gg 1$ then $|[f(\bm{A})]_{ij}|\ll 1$, which in turn means that depending on the sparsity pattern of $\bm{A}$, there may only be few large entries in $f(\bm{A})$ with the rest being negligible. With this observation we can aim to recover the dominating entries of each row/column of $f(\bm{A})$. We solve this problem using compressed sensing by interpreting small entries as noise and the large entries as signals. [^3] Another possibility would be to identify the nonzero structure using the graph structure $(V,E)$ and $f$. This can quickly become intractable, and in some cases $f(\bm{A})$ can be sparse despite the graph structure suggesting otherwise, so here we opt for a generic approach.
Compressed sensing CandesRombergTao06; Donoho06 is a technique in signal processing for efficiently reconstructing a sparse signal. Let $\bm{Y}\in\mathbb{R}^{n\times s}$ be a so-called sensing operator and $\bm{v}\in\mathbb{R}^{n}$ be an unknown sparse vector. Then the problem of recovering $\bm{v}$ using the measurements $\bm{y}=\bm{Y}^{T}\bm{v}$ is a combinatorial optimization problem given by
@@
where $\left\lVert\bm{z}\right\rVert_{0}$ denotes the number of nonzeros in the vector $\bm{z}$. There are two established classes of algorithms, among others, to solve (2) in the compressed sensing literature, namely basis pursuit or $\ell_{1}$-minimization using convex optimization CandesTao05, and greedy methods BlumensathDaviesRiling12. For large $n$, convex optimization can become infeasible (especially given that we need to solve (2) $n$ times to recover the whole matrix), so we use the cheaper option, greedy methods. Greedy methods are iterative and at each iteration attempt to find the correct location and the values of the dominant entries.
Let $\bm{v}\in\mathbb{R}^{n}$ be a $k$-sparse vector, that is, a vector with at most $k$ nonzero entries. To recover $\bm{v}$ with theoretical guarantee we typically require at least $s=\mathcal{O}(k\log n)$ measurements. In practice, $s=\mathcal{O}(k)$ measurements seem to work well BlanchardTanner15. Each measurement is taken through vector-vector multiply with $\bm{v}$. The three classes of frequently used sensing operators $\bm{Y}$ are: a Gaussian matrix $\bm{\mathcal{N}}$ with entries drawn from i.i.d. $\mathcal{N}(0,s^{-1})$, $s$ columns of the discrete cosine transform matrix $\bm{\mathcal{C}}$ sampled uniformly at random and a sparse matrix $\bm{\mathcal{S}}_{\xi}$ with $\xi$ nonzero entries per row with the support set drawn uniformly at random and the nonzero values drawn uniformly at random from $\pm\xi^{-1/2}$. Here, $\xi$ is the sparsity parameter for the sparse sensing operator.
There are a number of greedy methods available, including Normalized Iterative Hard Thresholding (NIHT) NIHT, Hard Thresholding Pursuit (HTP) HTP and Compressive Sampling Matching Pursuit (CoSaMP) CoSaMP. Among the key differences and the trade-offs between the different methods is the cost per iteration in order to speed up support detection and the number of iterations for better asymptotic convergence rate. Many greedy methods have uniform recovery guarantees if the sensing operator $\bm{Y}$ has sufficiently small restricted isometry constants csbook. [^4] Under mild assumptions on $s$, the three classes of sensing operators described in the previous paragraph have small restricted isometry constants with high probability. This means that if we want to recover a matrix then we can use the same sensing operator $\bm{Y}$ for the recovery of all rows of a matrix, which makes it much more efficient than recovering $n$ rows of a matrix individually using different sensing operators.
@@
Here $H_{k}$ is the hard thresholding operator that sends all but the largest $k$ entries (in absolute value) of the input vector to 0. The stepsize $\mu$ in NIHT is chosen to be optimal when the support of the current iterate is the same as the sparsest solution. Assuming that $\bm{Y}$ satisfies the restricted isometry property (RIP), then at iteration $N$, the output $\bm{v}^{(N)}$ of NIHT satisfies
@@
Further details about NIHT can be found in NIHT.
The dominating cost for the greedy methods is the matrix-vector product between the sensing operator or the transpose of the sensing operator and the unknown sparse signal, i.e., the cost of $\bm{v}\mapsto\bm{Y}\bm{v}$ and $\bm{w}\mapsto\bm{Y}^{T}\bm{w}$ respectively. If the sensing operator is a Gaussian $\bm{\mathcal{N}}$ then the cost is $\mathcal{O}(ns)$, if $\bm{Y}$ is a subsampled discrete cosine matrix $\bm{\mathcal{C}}$ then the cost is $\mathcal{O}(n\log n)$ or $\mathcal{O}(n\log s)$ if we use the subsampled FFT algorithm rokhlintygert2008 and lastly, if $\bm{Y}$ is a sparse matrix $\bm{\mathcal{S}}_{\xi}$ then the cost is $\mathcal{O}(\xi n)$.
@@ 2.1 Strategy for recovering general sparse matrices
where $C_{B}>0$ and $0<\lambda_{B}<1$ are constants and $d(i,j)\in\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is some bivariate function that captures the magnitude of the entries of $\bm{B}$. Functions of sparse matrices $\bm{B}=f(\bm{A})$ with a smooth $f$ is a special case of $\bm{B}$ that satisfies (1). Another special case is $\bm{B}=\bm{A}$, which recovers the sparse matrix $\bm{A}$ itself. The zero entries of $\bm{B}$ would correspond to $d(i,j)=\infty$ in (5).
We are concerned with recovering the whole matrix. We approximate the dominant entries of each row of $\bm{B}$ by taking measurements through matrix-vector products $\bm{B}\bm{y}_{i}$ where $\bm{y}_{i}$ is the $i$th column of the sensing operator $\bm{Y}\in\mathbb{R}^{n\times s}$. [^6] We can also recover the dominant entries of $\bm{B}$ column by column using $\bm{B}^{T}\bm{Y}$ if the matrix-vector product $\bm{x}\mapsto\bm{B}^{T}\bm{x}$ is available. Let $\texttt{CS}(\bm{Y},\bm{b}_{i}\bm{Y},k)$ be a function that solves (2) where $\bm{Y}\in\mathbb{R}^{n\times s}$ is the sensing operator, $\bm{b}_{i}\bm{Y}\in\mathbb{R}^{1\times s}$ are the measurements where $\bm{b}_{i}$ is the $i$th row of $\bm{B}$ (i.e., $\bm{b}_{i}\bm{Y}$ is the $i$th row of $\bm{BY}$) and $k$ is the sparsity parameter. For the compressed sensing algorithm CS, we can take, for example $\texttt{CS}=$ NIHT. The algorithm to recover the dominant entries of each row of $\bm{B}$ is given below in Algorithm 1, which we call SpaMRAM (Sparse Matrix Recovery Algorithm using Matvecs).
@@
[…]
3:Draw a sensing operator $\bm{Y}\in\mathbb{R}^{n\times s}$, e.g. Gaussian. $\triangleright$ rec. $s=\lceil 2k\log(n/k)\rceil$
@@
5:for i=1,2,,ni=1,2,...,n do
6:  $\widehat{\bm{B}}(i,:)=\texttt{CS}(\bm{Y},\bm{F}_{s}(i,:),k)$ $\triangleright$ Using compressed sensing to approximate the rows of $\bm{B}$
@@
The algorithm SpaMRAM is a straightforward application of compressed sensing techniques used to recover a sparse matrix instead of a vector, particularly when the goal is to recover $\bm{A}$ itself. We are nonetheless unaware of existing work that spells out explicitly, and SpaMRAM can be significantly more efficient than alternative approaches HermanStrohmer2009; PfranderRauhutTanner2008 that vectorize the matrix. Since most sensing matrices satisfy the RIP with failure probability exponentially decaying in $s$ csbook, by the union bound we can conclude that the whole matrix can be recovered with high probability.
@@
When $\bm{B}=f(\bm{A})$ for sparse $\bm{A}$, we can evaluate $f(\bm{A})\bm{Y}$ using Krylov methods. Assuming that $f$ is analytic in a region containing the eigenvalues of $\bm{A}$, [^7] It also suffices for $f$ to be approximable to high accuracy with a low-degree polynomial. the Krylov method converges (to sufficient accuracy) in $\mathcal{O}(1)$ iterations, then the complexity of SpaMRAM is $\mathcal{O}(s\cdot nnz(\bm{A})+n^{2}\log n)=\mathcal{O}(k\log n\cdot nnz(\bm{A})+n^{2}\log n)$ which consists of $\mathcal{O}(s\cdot nnz(\bm{A}))=\mathcal{O}(k\log n\cdot nnz(\bm{A}))$ flops for computing $s=\mathcal{O}(k\log n)$ matrix-vector products with $f(\bm{A})$ and $\mathcal{O}(n^{2}\log n)$ flops for solving $n$ compressed sensing problems. Here $nnz(\bm{A})$ is the number of nonzero entries of $\bm{A}$.
@@ 2.1.2 Dependency of the compressed sensing solver in SpaMRAM
The majority of the computational task in SpaMRAM (Algorithm 1) relies on compressed sensing. There are three main parameter choices in relation to compressed sensing for SpaMRAM: the sensing operator $\bm{Y}$, the number of measurements $s$ and the compressed sensing solver CS. For the solver, we suggest NIHT as it is the simplest and has a low cost per iteration. For better asymptotic guarantee, but with a higher cost per iteration, we suggest CoSaMP or its variants, e.g. CSMPSP discussed in BlanchardTanner15. For the sensing operator $\bm{Y}$, all three choices mentioned previously are good choices. For robustness, we recommend the Gaussian matrix $\bm{\mathcal{N}}$ and for speed, we recommend the sparse matrix $\bm{\mathcal{S}}_{\xi}$ with the sparsity parameter $\xi=\min\{8,s\}$ if a good sparse matrix library is available, otherwise we recommend the subsampled DCT matrix $\bm{\mathcal{C}}$. Lastly, for the number of measurements $s$ we recommend $s=\lceil 2k\log\left(\frac{n}{k}\right)\rceil$ for theoretical guarantee Donoho06; CandesRombergTao06 as suggested in SpaMRAM. However, a constant multiple of the sparsity parameter $k$, say $s=10k$, is observed to work well in practice BlanchardTanner15; csbook.
@@ 2.1.3 A posteriori error estimate for SpaMRAM
The behaviour of SpaMRAM can be difficult to grasp as it relies on various parameters that are sometimes implicit, e.g., (5). Therefore, we can incorporate a posteriori error testing for SpaMRAM to observe how well it might perform on a particular problem. Let us try to estimate the error using known quantities. We have the sensing operator $\bm{Y}$, the measurements $\bm{F}_{s}\in\mathbb{R}^{n\times s}$ and the output of SpaMRAM, $\hat{\bm{B}}$. If the matrix-vector products (line $2$ of SpaMRAM) have been computed exactly then we have $\bm{B}\bm{Y}=\bm{F}_{s}$, so we try to estimate the error by determining how close the measurements computed using the output of SpaMRAM, i.e. $\hat{\bm{B}}\bm{Y}$, are to the computed measurements $\bm{F}_{s}$. We define
@@
as the relative error of the computed solution of SpaMRAM quantified through the measurements. Note that the error estimate $\delta_{RE}^{(S)}$ can be computed a posteriori in $\mathcal{O}(nks)$ as $\hat{\bm{B}}$ has at most $k$ nonzero entries in each row. The error estimate $\delta_{RE}^{(S)}$ can be used as a heuristic for how SpaMRAM performs. The main motivation for this approach is that a small number of Gaussian measurements estimate the norm well GrattonTitley-Peloquin2018; mt20. In our case, we essentially have $\bm{F}_{s}=\bm{B}\bm{Y}$ and $\hat{\bm{B}}\bm{Y}$ giving us $(\bm{B}-\hat{\bm{B}})\bm{Y}$, which is a small number of Gaussian measurements of the absolute error $\bm{B}-\hat{\bm{B}}$ if $\bm{Y}$ is the Gaussian sensing operator. [^8] For theoretical guarantee, $\bm{Y}$ needs to be independent from $\bm{B}$ and $\hat{\bm{B}}$ GrattonTitley-Peloquin2018, however the output of SpaMRAM, $\hat{\bm{B}}$ is dependent on $\bm{Y}$. Nonetheless, the dependency is rather weak so $\delta_{RE}^{(S)}$ should be a good error estimate in practice. For theoretical guarantee, one can take, say $5$ extra measurements for error estimation. Therefore $\delta_{RE}$ provides a good guidance for how small the error is for SpaMRAM. For the other sensing operators, the behaviour is similar mt20. Later, we show in a numerical experiment (Figure 2) in Section 4 that $\delta_{RE}^{(S)}$ can estimate the performance of SpaMRAM well.